the lenght of a rectangular garden is 5 meters more than twice the width. the perimeter is 34 meters. find the lenght and width of the garden
Answer by checkley77(12844)

L=2W+5
2L+2W=P
2(2W+5)+2W=34
4W+10+2W=34
6W=34-10
6W=24
W=24/6
W=4 METERS ANS. FOR THE WIDTH.
L=2*4+5=8+5=13 METERS ANS. FOR THE LENGTH.
PROOF:
2*13+2*4=34
26+8=34
34=34